OpenStudy (anonymous):

How do I determine x for this pentagon? The angles are x, 1.5x, 1.5x, x,x

OpenStudy (anonymous):

first find the sum of interior angles of a pentagon by using the formula (n-2) * 180

OpenStudy (anonymous):

oh ok ill try that

OpenStudy (anonymous):

would the anwser be -90?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

(5-2) * 180 = 3 * 180 = 540 x + 1.5x + 1.5x + x +x = 540

OpenStudy (anonymous):

collect like terms and solve for x

OpenStudy (anonymous):

i forgot to add 1.5 and 1.5 together... T-T

OpenStudy (anonymous):

6x = 540 x = 90

OpenStudy (anonymous):

i asked my father and he doesnt even understand... i dont quite understand...

OpenStudy (anonymous):

Which part do you not understand?

OpenStudy (anonymous):

well... all of it...

OpenStudy (anonymous):

What is the formula for finding the sum of all interior angles for an n sided shape?

OpenStudy (anonymous):

s=(n-2)180

OpenStudy (anonymous):

Good! so what is the sum of all interior angles for a pentagon, which is a 5-side shape?

OpenStudy (anonymous):

thats where i am a little hasey

OpenStudy (anonymous):

Let's do this slowly. Plug in 5 as n

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then multiply it by 180 and you get 540?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

HAHA FINALLY!!!

OpenStudy (anonymous):

Given that Tthe angles are x, 1.5x, 1.5x, x,x

OpenStudy (anonymous):

sum of these angles is also equal to 540

OpenStudy (anonymous):

so all of the angles added together = 540?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i drew an image of the shape would you want to see?

OpenStudy (anonymous):

sure, but it's not required

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then i wont

OpenStudy (anonymous):

ok so do you understand it much better?

OpenStudy (anonymous):

sort of.. in class my teacher saud something about subtracting the sum with 180

OpenStudy (anonymous):

subtracting 180 is for triangles, which is a 3-sided shape

OpenStudy (anonymous):

oh ok... thank you for taking your time and explaining this to me...

OpenStudy (anonymous):

no problem